Database last updated on Tue, 27 Jan 2026 15:40:51 +0000 by localhost
Critical warnings from b-files' comments for days between 2025110 and 2025110 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 109 05 33 37 Motor #2: appears stuck with a reference sensor activated. | 2 | 1 | B11025.224 | TORON | #224 | 05:27:21 | | 1 | -1 |
| rl: | 2025 110 07 42 16 New value set for watermark SUPPLY.VOLTAGE.LOW | 12 | 3 | B11025.224 | TORON | #224 | 05:27:26 | | 1 | -1 |
Instrument constant changes for the week before 2025110 (1 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 4 rows)
HG burned out? (AP results when HG is ON) (0 sec, 1 rows)
Non-critical warnings from b-files' comments for days between 2025110 and 2025110 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 109 05 31 08 Reset Warm -> Tepid from 2025 109 05 31 03 until 2025 109 05 31 03 . Counts: 1,0 | 1 | 3 | B11025.224 | TORON | #224 | 05:27:18 | | 1 | -1 |
| ds: | Suppressed oscillation of filter wheel. | 1 | 5 | B11025.224 | TORON | #224 | 12:25:45 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | ds | dt | ed | fm | fz | hg | rs | sl | uv | zs |
| 110 | TORONTO | #224 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.143) | 81 | 2 | 2 | 3 | 9 | 22 | 2 | 7 | 29 | 28 |
(0 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 224 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 224 | 2025111 | 0-1 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 224 | 14 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #224 | 2670440 | 2670720 | 0 | yes |